The percent of N in 66% pure (NH402 4 sample is A-32,b-28

Grade:10

1 Answers

somi teez
6 years ago
i think your compound is (nh4)2 so4
now,
molecular mass of compound=2*(14+1*4)+32+16*4
                                           =2*18+32+64
                                          = 36+96
                                          =132g
atomic weight of nitrogen=28g
so,
  % of nitrogen in 66% =28/(66/100 *132) %=32 % approx